# Counterfactual Sensitivity in Quantitative Spatial Models

### v2

@@ D.2.1 Calibration Procedure
is an unbiased estimator. Towards estimating $\log F_{ij,t}$ and $s_{ij}^{2}$, we can combine the model equations to find:
@@
D.2.2 Computational Implementation Details
@@
When estimating the prior distribution of the true underlying trade flows, I use the distance dataset from Mayer and Zignago (2011). For the distance between countries and the “Rest of World”, I take the average of the distances to all other countries that are considered in Adao, Costinot, and Donaldson (2017).
@@
Furthermore, for the computational implementation of the bounds that incorporate both estimation error and measurement error, I use the code provided by the authors of Adao, Costinot, and Donaldson (2017) to account for estimation error. That code is somewhat unstable so I make two modifications. Firstly, for some draws of the structural parameter the code was not converging. I opted to ignore these draws when constructing the bounds. Secondly, for the adversarial bounds as in Remark 6 I formally need to take the derivative of the upper and lower bound functions that incorporate estimation error with respect to the observed trade flows, which are $\frac{\partial g_{L_{\hat{k}^{E}}^{\beta}}}{\partial D_{O}}|_{\tilde{D}_{O}}^{{}^{\prime}}$ and $\frac{\partial g_{U_{\hat{k}^{E}}^{\beta}}}{\partial D_{O}}|_{\tilde{D}_{O}}^{{}^{\prime}}$. Since the code that generates these bounds is unstable, I approximate these derivatives by the Jacobian of the point estimate with respect to the observed trade flows, $\frac{\partial g_{\hat{k}}}{\partial D_{O}}|_{\tilde{D}_{O}}^{{}^{\prime}}$.
D.3 Supplementary Analyses
@@
D.3.1 Flat Prior and Homoskedastic Lower Bound for Measurement Error Variances
@@
The first supplementary analysis uses a flat prior and a uniform measurement error variance for all trade flows. This measurement error variance is the average of $\left\{\mathring{\varsigma}_{ij}^{2}\right\}$ from Equation (16), but only considering the countries France, Germany, Great Britain, Japan and the United States, which are all generally considered to be reliable reporters of trade data. This average can hence be interpreted as a lower bound on the measurement error variance, and it equals $\varsigma_{LB}^{2}=0.0048$. The resulting confidence and credible intervals are plotted in Figure 2. Even in this conservative case, the credible set that incorporates both estimation error and measurement error still is large, and for the period 2002-2011 it becomes ambiguous whether joining the WTO had a positive or negative effect on China’s welfare.
D.3.2 Winsorized Measurement Error Variances
@@
The distribution of measurement error variances has a heavy right tail, with the noisiest bilateral trade flow the one from Mexico to Russia with a measurement error variance of 1.14. One might be worried that this heavy tail drives the sensitivity to mismeasurement. Figure 3 replicates Figure 1 but now winsorizing the measurement error variances at 0.2. There are no notable differences between Figures 3 and 1.
D.3.3 Using Adversarial Uncertainty Quantification
Figure 4 replicates Figure 1 but now using the adversarial approach rather than the EB approach. The confidence regions are generally comparable, with the adversarial confidence regions that combine estimation error and measurement error being less extreme than the EB version. For the adversarial approach I use the $10\%$ confidence interval of the measurement error distribution to choose $\boldsymbol{\epsilon}$. This amounts to choosing $\epsilon_{ij}=0.13\varsigma_{ij}$ for all $i,j$, where $\epsilon_{ij}$ is the adversarial bound corresponding to the log trade flow between country $i$ and $j$. Figure 5 depicts the sensitivity analysis plot for the confidence intervals of the measurement error for choosing $\boldsymbol{\epsilon}$, for the changes in welfare in 2011. Interestingly, the lower bounds for different years seems to converge, while the upper bounds diverge.
D.3.4 Probability of Change in Welfare in China Being Larger Than 1
@@
We can learn more from the posterior distributions than just credible sets. It might be of interest what the posterior probability is that the welfare in China increased as a result of the China shock. For the cases with only estimation error or measurement error, this is a trivial exercise. For the case with both estimation error and measurement error I use
$\frac{1}{B}\sum_{b=1}^{B}\min\left\{\max\left\{\frac{U_{b,\hat{W}_{\mathrm{China},t}^{E}}^{\beta}-1}{U_{b,\hat{W}_{\mathrm{China},t}^{E}}^{\beta}-L_{b,\hat{W}_{\mathrm{China},t}^{E}}^{\beta}},0\right\},1\right\}\left(1-\beta\right).$
@@
For the period 1996-2011, the results are plotted in Figure 8. We observe that this probability depends on whether we consider only estimation error, only measurement error, or both. For each of those three cases, the probability seems to be the highest in the period 2006-2007.
D.3.5 Testing Normality Assumption and Gravity Model for the Prior
@@
As outlined in Remark 3, we can check the reasonableness of the normality assumption on the prior by comparing the histogram of the normalized residuals with the probability density function of a standardized normal distribution. The result can be found in Figure 6. It follows that the normality assumption seems reasonable.
@@
Concerning the gravity model, for the year 2011 the regression for the prior mean in Equation (13) has an adjusted R-squared of $0.950$, and the coefficient on log distance is $-0.275$ with a t-statistic of $3.380$. Furthermore, Figure 7 follows Allen and Arkolakis (2018) by plotting a linear and nonparametric fit of log trade flows against log distance, after partitioning out the origin and destination fixed effects. Together, the high adjusted R-squared and the good performance of the linear fit, imply that the gravity model is a reasonable choice for this application.
D.3.6 Variance Decomposition
@@
One might wonder what drives uncertainty in this application, and I use a simple first-order variance decomposition to explore this. For example looking at the change in welfare for China in 2011 due to the China shock, its variance is for 47% explained by the ten most important bilateral trade flows (out of 22,032). Table 3 lists these ten bilateral trade flows.
@@
As expected, many of these important bilateral trade flows involve China and its most important trade partners from 2011. It is however remarkable that Australia and the United States show up multiple times, especially their trade flows with seemingly irrelevant countries for China’s welfare like Slovakia and the Baltic states in the years 1999 and 2000. This can be explained by the fact that in the estimation of the structural parameter, which is the elasticity of substitution, only the data from Australia and the United States are used. This shows that the elasticity of substitution is a strong driver of sensitivity, and that in its estimation some peculiar trade flows play an important role.
Appendix E Details for Application Allen and Arkolakis (2022)
E.1 Model Details
@@
$\hat{y}_{i}^{\frac{1+\gamma\delta+\gamma}{1+\gamma\delta}}\hat{\ell}_{i}^{\frac{-\theta\left(1+\alpha+\gamma\delta\left(\alpha+\beta\right)\right)}{1+\gamma\delta}}$ $=\hat{\chi}\left(\frac{y_{i}\bar{Y}}{y_{i}\bar{Y}+\sum_{k}F_{ik}}\right)\hat{y}_{i}^{\frac{1+\gamma\delta+\gamma}{1+\gamma\delta}}\hat{\ell}_{i}^{\frac{\gamma\left(\beta-1\right)}{1+\gamma\delta}}$
$\qquad+\sum_{j}\left(\frac{F_{ij}}{y_{i}\bar{Y}+\sum_{k}F_{ik}}\right)\hat{\tau}_{ij}^{\frac{-\gamma}{1+\gamma\delta}}\hat{y}_{j}^{\frac{1+\gamma}{1+\gamma\delta}}\hat{\ell}_{j}^{\frac{-\gamma\left(1+\alpha\right)}{1+\gamma\delta}}$
$\hat{y}_{i}^{\frac{-\gamma\left(1-\delta\right)}{1+\gamma\delta}}\hat{\ell}_{i}^{\frac{\gamma\left(1-\beta-\gamma\delta\left(\alpha+\beta\right)\right)}{1+\gamma\delta}}$ $=\hat{\chi}\left(\frac{y_{i}\bar{Y}}{y_{i}\bar{Y}+\sum_{k}F_{ki}}\right)\hat{y}_{i}^{\frac{-\gamma\left(1-\delta\right)}{1+\gamma\delta}}\hat{\ell}_{i}^{\frac{\gamma\left(\alpha+1\right)}{1+\gamma\delta}}$
$\qquad+\sum_{j}\left(\frac{F_{ji}}{y_{i}\bar{Y}+\sum_{k}F_{ki}}\right)\hat{\tau}_{ij}^{\frac{-\gamma}{1+\gamma\delta}}\hat{y}_{j}^{\frac{-\gamma}{1+\gamma\delta}}\hat{\ell}_{j}^{\frac{\gamma\left(1-\beta\right)}{1+\gamma\delta}}.$
@@
$\hat{W}^{E}=\frac{\hat{\chi}^{1/\gamma}}{\bar{L}^{\alpha+\beta}}.$
E.2 Estimation details
E.2.1 Calibration Procedure
The calibration procedure is similar to the one for the application from Adao, Costinot, and Donaldson (2017). Combining the prior and the measurement error model now leads to the following hierarchical model:
@@ E.2.1 Calibration Procedure
$\hat{s}^{2}$ $=\widehat{\operatorname{Var}}\left(\log\tilde{F}_{ij}-\hat{\beta}\log\mathrm{dist}_{ij}+\hat{\alpha}_{i}^{\mathrm{orig}}+\hat{\alpha}_{j}^{\mathrm{dest}}\right)-0.05.$
E.2.2 Computational Implementation Details
@@
When I run the code from Allen and Arkolakis (2022), the returns of investment for the links systematically differ slightly from the ones in the paper. I scale my estimates so that the unperturbed estimates align with the ones in the paper. For the procedure that combines measurement error and estimation error, I use the delta method to obtain the frequentist confidence interval that accounts for estimation error rather than a bootstrap procedure for computational reasons.
For the adversarial approach, I only consider measurement error of traffic flows that are adjacent. This is mainly for computational reasons, because calculating the numerical derivative of the returns on investment with respect to a traffic flow is quite computationally expensive. In simulations I observe that these adjacent traffic flows capture most of the variation, in that including the non-adjacent traffic flows will only slightly widen the confidence bounds. Also, for the adversarial approach, similar as for Adao, Costinot, and Donaldson (2017), I approximate the Jacobian at the upper and lower bound by the Jacobian of the point estimate with respect to the traffic flows.
E.3 Supplementary Analyses
E.3.1 Using Adversarial Uncertainty Quantification
The adversarial bounds for the levels and differences of the top three links can be found in Tables 4 and 5. The adversarial approach uses the $10\%$ confidence interval of the measurement error distribution to choose $\epsilon$. The resulting confidence regions are small compared to the ones obtained using the EB approach. Since in Allen and Arkolakis (2022) no estimation error is considered, in Figure 9 I construct the sensitivity analysis plot for the confidence intervals under the assumption of no estimation error. It turns out that in this case the ranking does not change even for large values of the adversarial parameter $\epsilon$.
E.3.2 Probability that Rankings are Reversed
@@
We can learn more from the posterior distributions than just credible sets. It might be of interest what the posterior probability is that the ranking of the three links are reversed. When we consider only measurement error, this probability that the ranking between link 1 and link 2 is reversed equals 0.000 and that the ranking between link 2 and link 3 is reversed equals 0.023. When we consider only estimation error these probabilities change to 0.402 and 0.236 respectively. When we consider both measurement error and estimation error simultaneously, which is calculated using the same approach as in SectionD.3.4, these increase further to 0.600 and 0.584, respectively.
E.3.3 Testing Normality Assumption and Gravity Model for the Prior
@@
We can again check the reasonableness of the normality assumption on the prior as per Remark 3. The result can be found in Figure 10, and it follows that the normality assumption is less reasonable then as for application of Adao, Costinot, and Donaldson (2017).
@@
Concerning the gravity model, the regression for the prior mean in Equation (12) has an adjusted R-squared of $0.9995$, and the coefficient on log distance is $1.003$ with a t-statistic of $1138$. It follows that log distance is an important driver of log traffic flows, but not in a negative way as is common in gravity models. Furthermore, Figure 11 follows Allen and Arkolakis (2018) by plotting a linear and nonparametric fit of log traffic flows against log distance, after partitioning out the origin and destination fixed effects. Together, the high adjusted R-squared and the good performance of the linear fit, imply that the gravity model is a reasonable choice for this application.
Appendix F Misspecification of the Measurement Error Model
@@
Let $L\left(x\right)\propto p\left(\tilde{x}|x\right)$ denote the likelihood function of the noisy data $\left(\tilde{X}_{O},\tilde{N}_{O}\right)$ given the true data $\left(X_{O},N_{O}\right)$. For a given $c>1$ define a density-ratio class of distributions to be the set of all conditional distributions for $\left(\tilde{X}_{O},\tilde{N}_{O}\right)$ with pdf $p$ such that
$p\in\mathcal{R}_{c}=\left\{p\in\mathcal{D}:L\left(x\right)\leq p\left(x\right)\leq c\cdot L\left(x\right)\ \forall x\in\mathcal{X}_{O}\times\mathcal{N}_{O}\right\},$
@@
for $\mathcal{D}$ the set of all pdfs. For any function $g\left(\cdot\right)$, we are then interested in the bounds
$\overline{\mathbb{E}}_{\pi}\left[g\left(X_{O},N_{O}\right)|\tilde{X}_{O},\tilde{N}_{O}\right]$ $=\underset{p\in\mathcal{R}_{c}}{\sup}\mathbb{E}_{\pi}\left[g\left(X_{O},N_{O}\right)|\tilde{X}_{O},\tilde{N}_{O}\right]$
$\underline{\mathbb{E}}_{\pi}\left[g\left(X_{O},N_{O}\right)|\tilde{X}_{O},\tilde{N}_{O}\right]$ $=\underset{p\in\mathcal{R}_{c}}{\inf}\mathbb{E}_{\pi}\left[g\left(X_{O},N_{O}\right)|\tilde{X}_{O},\tilde{N}_{O}\right].$
@@
Denoting with $\pi\left(x\right)$ the pdf of the specified prior-distribution, from Proposition 1 in Geweke and Petrella (1998) we then have the following proposition:
@@
$\overline{\mathbb{E}}_{\pi}\left[g\left(X_{O},N_{O}\right)|\tilde{X}_{O},\tilde{N}_{O}\right]$ is the unique solution $t$ to
@@
and $\underline{\mathbb{E}}_{\pi}\left[g\left(X_{O},N_{O}\right)|\tilde{X}_{O},\tilde{N}_{O}\right]=-\overline{\mathbb{E}}_{\pi}\left[g\left(X_{O},N_{O}\right)|\tilde{X}_{O},\tilde{N}_{O}\right]$.
For the normal-normal model, this leads to the following corollary:
@@
In the special case that $L\left(x\right)\propto\phi\left(x;\mu_{m},\Sigma_{m}\right)$ and $\pi\left(x\right)\propto\phi\left(x;\mu_{p},\Sigma_{p}\right)$, $\overline{\mathbb{E}}_{\pi}\left[g\left(X_{O},N_{O}\right)|\tilde{X}_{O},\tilde{N}_{O}\right]$ is the unique solution $t$ to
@@
$\mu=\left(\Sigma_{m}^{-1}+\Sigma_{p}^{-1}\right)^{-1}\left(\Sigma_{m}^{-1}\mu_{m}+\Sigma_{p}^{-1}\mu_{p}\right)\quad\mathrm{and}\quad\Sigma=\left(\Sigma_{m}^{-1}+\Sigma_{p}^{-1}\right)^{-1}.$
We can apply these results for the posterior mean by choosing $g\left(X_{O},N_{O}\right)=g_{T}\left(X_{O},N_{O}\right)$, and we could target $q_{\alpha}$, the $\alpha$-th quantile of the posterior $T|\tilde{X}_{O},\tilde{N}_{O}$, by choosing
$g\left(X_{O},N_{O}\right)=\inf\left\{y\in\mathbb{R}:\mathbb{P}\left[g_{T}\left(X_{O},N_{O}\right)\leq y\right]\geq\alpha\right\}.$
Appendix G Proofs
Proof of Proposition 2
Consider the problem
$\boldsymbol{\eta}_{\max}=\arg\max_{\boldsymbol{\eta}:\left|\boldsymbol{\eta}_{j}\right|\leq\boldsymbol{\epsilon}_{j}\forall j}\tilde{T}_{\boldsymbol{\eta}}-\tilde{T}_{\boldsymbol{0}}.$
Note that if $\tilde{T}_{\boldsymbol{\eta}}$ is approximately linear in $\tilde{D}_{O}$ on $\left[\tilde{D}_{O}-\boldsymbol{\epsilon},\tilde{D}_{O}+\boldsymbol{\epsilon}\right]$, we have
$\tilde{T}_{\boldsymbol{\eta}}\approx\tilde{T}_{\boldsymbol{0}}+\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)\boldsymbol{\eta}.$
Hence, to maximize $\tilde{T}_{\boldsymbol{\eta}}-\tilde{T}_{\boldsymbol{0}}$ we will always choose $\boldsymbol{\eta}_{j}$ at the boundary of $\left[-\boldsymbol{\epsilon}_{j},\boldsymbol{\epsilon}_{j}\right]$, where the choice of which boundary is governed by the sign of the $j$-th element of $\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)$.
Appendix H Appendix Tables and Figures

### v3

@@ C.2.1 Calibration Procedure
is an unbiased estimator for $\varsigma_{ij}^{2}$. Towards estimating $\log F_{ij,t}$ and $s_{ij}^{2}$, we can combine the model equations to find:
@@
C.2.2 Computational Implementation Details
@@
When estimating the prior distribution of the true underlying trade flows, I use the distance dataset from Mayer and Zignago (2011). For the distance between countries and the “Rest of World”, I take the average of the distances to all other countries that are considered in Adao, Costinot, and Donaldson (2017). For a handful of bilateral trade flows, the estimated prior variance was negative. I opted to drop these particular trade flows.
@@
For the computational implementation of the bounds that incorporate both estimation error and measurement error, I use the code provided by the authors of Adao, Costinot, and Donaldson (2017) to account for estimation error. For some draws of the structural parameter the code was not converging. I opted to ignore these draws when constructing the bounds.
C.3 Supplementary Analyses
@@
C.3.1 Flat Prior and Homoskedastic Lower Bound for Measurement Error Variances
@@
The first supplementary analysis uses a flat prior distribution and a uniform measurement error variance for all trade flows. This measurement error variance is the average of $\left\{\mathring{\varsigma}_{ij}^{2}\right\}$ from Equation (13), but only considering the countries France, Germany, Great Britain, Japan and the United States, which are all generally considered to be reliable reporters of trade data. This average can hence be interpreted as a lower bound on the measurement error variance, and it equals $\varsigma_{LB}^{2}=0.0048$. The resulting intervals are plotted in Figure 2. Even in this conservative case, the interval that incorporates both estimation error and measurement error still is large, and for the period 2002-2011, with the exception of 2009, it becomes ambiguous whether joining the WTO had a positive or negative effect on China’s welfare.
C.3.2 Winsorized Measurement Error Variances
@@
The distribution of measurement error variances has a heavy right tail, with the noisiest bilateral trade flow the one from Mexico to Russia with a measurement error variance of 1.14. One might be worried that this heavy tail drives the sensitivity to mismeasurement. Figure 3 replicates Figure 1 but now winsorizing the measurement error variances at 0.1. There are no notable differences between Figures 3 and 1.
C.3.3 Using $\mathcal{C}^{2}$ instead of $\mathcal{C}^{1}$
Following the discussion in 3.3, Figure 4 plots the interval $\mathcal{C}^{2}$ for the change in China’s welfare. Indeed, the interval that combines estimation and measurement error becomes extremely wide.
C.3.4 Probability of Change in Welfare in China Being Larger Than 1
@@
We can learn more from the posterior distributions than just intervals. It might be of interest what the expected probability is that the welfare in China increased as a result of the China shock:
$\mathbb{E}_{\pi^{ME}}\left[Pr_{\pi^{EE}}\left\{\hat{W}_{\mathrm{China},t}>1|\left\{\log F_{ij,t}\right\}\right\}|\left\{\log\tilde{F}_{ij,t}\right\}\right].$
@@
For the period 1996-2011, these expected probabilities are plotted in Figure 5. We observe that this expected probability does not seem to depend on whether we consider only estimation error, only measurement error, or both. For each of those three cases, the expected probability seems to be the highest in the period 2006-2007.
C.3.5 Testing Normality Assumption and Gravity Model for the Prior
@@
As outlined in Remark 2, we can check the reasonableness of the normality assumption by comparing the histogram of the normalized residuals with the probability density function of a standardized normal distribution. The result can be found in Figure 6. It follows that the normality assumption seems reasonable.
@@
Concerning the gravity model, for the year 2011 the regression for the prior mean in Equation (10) has an adjusted R-squared of $0.950$, and the coefficient on log distance is $-0.275$ with a t-statistic of $3.380$. Furthermore, Figure 7 follows Allen and Arkolakis (2018) by plotting a linear and nonparametric fit of log trade flows against log distance, after partitioning out the origin and destination fixed effects. Together, the high adjusted R-squared and the good performance of the linear fit imply that the gravity model is a reasonable choice for this setting.
C.3.6 Variance Decomposition
@@
One might wonder what drives uncertainty in this setting, and I use a simple first-order variance decomposition to explore this. For example looking at the change in welfare for China in 2011 due to the China shock, its variance is for 47% explained by the ten most important bilateral trade flows (out of 22,032). Table 3 lists these ten bilateral trade flows.
@@
As expected, many of these important bilateral trade flows involve China and its most important trade partners from 2011. It is however remarkable that Australia and the United States show up multiple times, especially their trade flows with seemingly irrelevant countries for China’s welfare like Slovakia and the Baltic states in the years 1999 and 2000. This can be explained by the fact that in the estimation of the structural parameter, which is the elasticity of substitution, only the data from Australia and the United States are used.
Appendix D Details for Application Allen and Arkolakis (2022)
D.1 Model Details
@@
$\hat{y}_{i}^{\frac{1+\delta+\gamma}{1+\delta}}\hat{\ell}_{i}^{\frac{-\theta\left(1+\alpha+\delta\left(\alpha+\beta\right)\right)}{1+\delta}}$ $=\hat{\chi}\left(\frac{y_{i}\bar{Y}}{y_{i}\bar{Y}+\sum_{k}F_{ik}}\right)\hat{y}_{i}^{\frac{1+\delta+\gamma}{1+\delta}}\hat{\ell}_{i}^{\frac{\gamma\left(\beta-1\right)}{1+\delta}}$
$\qquad+\sum_{j}\left(\frac{F_{ij}}{y_{i}\bar{Y}+\sum_{k}F_{ik}}\right)\hat{\tau}_{ij}^{\frac{-\gamma}{1+\delta}}\hat{y}_{j}^{\frac{1+\gamma}{1+\delta}}\hat{\ell}_{j}^{\frac{-\gamma\left(1+\alpha\right)}{1+\delta}}$
$\hat{y}_{i}^{\frac{-\gamma+\delta}{1+\delta}}\hat{\ell}_{i}^{\frac{\gamma\left(1-\beta-\delta\left(\alpha+\beta\right)\right)}{1+\delta}}$ $=\hat{\chi}\left(\frac{y_{i}\bar{Y}}{y_{i}\bar{Y}+\sum_{k}F_{ki}}\right)\hat{y}_{i}^{\frac{-\gamma+\delta}{1+\delta}}\hat{\ell}_{i}^{\frac{\gamma\left(\alpha+1\right)}{1+\delta}}$
$\qquad+\sum_{j}\left(\frac{F_{ji}}{y_{i}\bar{Y}+\sum_{k}F_{ki}}\right)\hat{\tau}_{ij}^{\frac{-\gamma}{1+\delta}}\hat{y}_{j}^{\frac{-\gamma}{1+\delta}}\hat{\ell}_{j}^{\frac{\gamma\left(1-\beta\right)}{1+\delta}}.$
@@
$\hat{W}=\frac{\hat{\chi}^{1/\gamma}}{\bar{L}^{\alpha+\beta}}.$
D.2 Estimation details
D.2.1 Calibration Procedure
The calibration procedure is similar to the one for the setting from Adao, Costinot, and Donaldson (2017). Combining the prior distribution and the measurement error model now leads to the following hierarchical model:
@@ D.2.1 Calibration Procedure
$\hat{s}^{2}$ $=\widehat{\operatorname{Var}}\left(\log\tilde{F}_{ij}-\hat{\beta}\log\mathrm{dist}_{ij}-\hat{\alpha}_{i}^{\mathrm{orig}}-\hat{\alpha}_{j}^{\mathrm{dest}}\right)-0.05.$
D.2.2 Computational Implementation Details
@@
When I run the code from Allen and Arkolakis (2022), the returns of investment for the links systematically differ slightly from the ones in the paper. I scale my estimates so that the unperturbed estimates align with the ones in the paper.
D.3 Supplementary Analyses
D.3.1 Probability that Rankings are Reversed
@@
We can learn more from the posterior distributions than just intervals. It might be of interest what the expected probability is that the ranking of the three links are reversed. When we consider only estimation error, this expected probability that the ranking between link 1 and link 2 is reversed and the expected probability that the ranking between link 2 and link 3 is reversed both equal 0.000.When we consider only measurement error these expected probabilities change to 0.000 and 0.030 respectively. When we consider both measurement error and estimation error simultaneously, the expected probabilities equal 0.000 and 0.022, respectively.
D.3.2 Testing Normality Assumption and Gravity Model for the Prior
@@
We can again check the reasonableness of the normality assumption as per Remark 2. The result can be found in Figure 8, and it follows that the normality assumption is less reasonable compared to the setting of Adao, Costinot, and Donaldson (2017).
@@
Concerning the gravity model, the regression for the prior mean in Equation (9) has an adjusted R-squared of $0.9995$, and the coefficient on log distance is $1.003$ with a t-statistic of $1138$. It follows that log distance is an important driver of log traffic flows, but not in a negative way as is common in gravity models. Furthermore, Figure 9 follows Allen and Arkolakis (2018) by plotting a linear and nonparametric fit of log traffic flows against log distance, after partitioning out the origin and destination fixed effects. Together, the high adjusted R-squared and the good performance of the linear fit imply that the gravity model is a reasonable choice for this setting.
Appendix E Misspecification of the Measurement Error Model
We are interested in the potential effects of misspecification of the measurement error model. Specifically, focusing on the widely applicable default approach from Section 3.4, we would like to know how the quantiles of the posterior distribution of the counterfactual change variable of interest given the noisy flows change when the assumption of a normal measurement error model does not hold. Suppose for exposition that the structural parameter $\theta$ is known, so that we can obtain the posterior distribution $\pi\left(\hat{k}|\left\{\log\tilde{F}_{ij}\right\}\right)$.
@@ Appendix E Misspecification of the Measurement Error Model
Let $L\left(\left\{\log F_{ij}\right\}\right)\propto\pi\left(\left\{\log\tilde{F}_{ij}\right\}|\left\{\log F_{ij}\right\}\right)$ denote the likelihood function of the noisy log flows $\left\{\log\tilde{F}_{ij}\right\}$ given the true log flows $\left\{\log F_{ij}\right\}$. For a given $c>1$ define a density-ratio class of distributions to be the set of all conditional distributions for $\left\{\log\tilde{F}_{ij}\right\}$ with pdf $p$ such that
$p\in\mathcal{R}_{c}=\left\{p\in\mathcal{D}:L\left(x\right)\leq p\left(x\right)\leq c\cdot L\left(x\right)\ \forall x\in\mathbb{R}_{++}^{n\left(n+1\right)}\right\},$
@@
for $\mathcal{D}$ the set of all pdfs. For a generic function $h\left(\cdot\right)$, we are then interested in the bounds
$\overline{\mathbb{E}}_{\pi}\left[h\left(\left\{\log F_{ij}\right\}\right)|\left\{\log\tilde{F}_{ij}\right\}\right]$ $=\underset{p\in\mathcal{R}_{c}}{\sup}\mathbb{E}_{\pi}\left[h\left(\left\{\log F_{ij}\right\}\right)|\left\{\log\tilde{F}_{ij}\right\}\right]$
$\underline{\mathbb{E}}_{\pi}\left[h\left(\left\{\log F_{ij}\right\}\right)|\left\{\log\tilde{F}_{ij}\right\}\right]$ $=\underset{p\in\mathcal{R}_{c}}{\inf}\mathbb{E}_{\pi}\left[h\left(\left\{\log F_{ij}\right\}\right)|\left\{\log\tilde{F}_{ij}\right\}\right].$
@@
Denoting with $\pi\left(\left\{\log F_{ij}\right\}\right)$ the pdf of the specified prior distribution, from Proposition 1 in Geweke and Petrella (1998) we then have the following result:
@@
$\overline{\mathbb{E}}_{\pi}\left[h\left(\left\{\log F_{ij}\right\}\right)|\left\{\log\tilde{F}_{ij}\right\}\right]$ is the unique solution $t$ to
@@
and $\underline{\mathbb{E}}_{\pi}\left[h\left(\left\{\log F_{ij}\right\}\right)|\left\{\log\tilde{F}_{ij}\right\}\right]=-\overline{\mathbb{E}}_{\pi}\left[h\left(\left\{\log F_{ij}\right\}\right)|\left\{\log\tilde{F}_{ij}\right\}\right]$.
For the normal-normal model, denoting with $\phi\left(x;\mu,\Sigma\right)$ a multivariate normal probability density function with mean $\mu$ and variance-covariance matrix $\Sigma$, this leads to the following corollary:
@@
If $L\left(\left\{\log F_{ij}\right\}\right)\propto\phi\left(\left\{\log F_{ij}\right\};\mu_{m},\Sigma_{m}\right)$ and $\pi\left(\left\{\log F_{ij}\right\}\right)\propto\phi\left(\left\{\log F_{ij}\right\};\mu_{p},\Sigma_{p}\right)$, then $\overline{\mathbb{E}}_{\pi}\left[h\left(\left\{\log F_{ij}\right\}\right)|\left\{\log\tilde{F}_{ij}\right\}\right]$ is the unique solution $t$ to
@@
$\mu^{*}=\left(\Sigma_{m}^{-1}+\Sigma_{p}^{-1}\right)^{-1}\left(\Sigma_{m}^{-1}\mu_{m}+\Sigma_{p}^{-1}\mu_{p}\right)\quad\mathrm{and}\quad\Sigma^{*}=\left(\Sigma_{m}^{-1}+\Sigma_{p}^{-1}\right)^{-1}.$
In Assumption 2, we chose
$\mu_{m}$ =$=\left\{\log\tilde{F}_{ij}\right\}$
$\Sigma_{m}$ $=\varsigma^{2}\cdot I_{n\left(n-1\right)}$
$\mu_{p}$ =$=\left\{\beta\log\mathrm{dist}_{ij}+\alpha_{i}^{\mathrm{orig}}+\alpha_{j}^{\mathrm{dest}}\right\}$
$\Sigma_{p}$ $=s^{2}\cdot I_{n\left(n-1\right)},$
so that
$\mu^{*}$ =$=\left\{\frac{\varsigma^{2}}{\varsigma^{2}+s^{2}}\log\tilde{F}_{ij}+\frac{s^{2}}{\varsigma^{2}+s^{2}}\left(\beta\log\mathrm{dist}_{ij}+\alpha_{i}^{\mathrm{orig}}+\alpha_{j}^{\mathrm{dest}}\right)\right\}$
$\Sigma^{*}$ $=\frac{\varsigma^{2}s^{2}}{\varsigma^{2}+s^{2}}\cdot I_{n\left(n-1\right).}$
We can apply these results for the posterior mean $\mathbb{E}_{\pi}\left[\hat{k}|\left\{\log\tilde{F}_{ij}\right\}\right]$ by choosing $h\left(\left\{\log F_{ij}\right\}\right)=g_{\hat{k}}\left(\left\{\log F_{ij}\right\};\theta\right)$, and we could target $q_{\alpha}$, the $\alpha$-th quantile of the posterior distribution $\pi\left(\hat{k}|\left\{\log\tilde{F}_{ij}\right\}\right)$, by choosing
$h\left(\left\{\log F_{ij}\right\}\right)=\inf\left\{y\in\mathbb{R}:\mathbb{P}\left[g_{\hat{k}}\left(\left\{\log F_{ij}\right\};\theta\right)\leq y\right]\geq\alpha\right\}.$
Appendix F Appendix Tables and Figures